# Egor Cherepanov1,2Nikita Kachaev1Alexey K. Kovalev1,2Aleksandr I. Panov1,2 1AIRI, Moscow, Russia  2MIPT, Dolgoprudny, Russia

### v1

@@ Abstract
Memory is crucial for enabling agents to tackle complex tasks with temporal and spatial dependencies. While many reinforcement learning (RL) algorithms incorporate memory, the field lacks a universal benchmark to assess an agent’s memory capabilities across diverse scenarios. This gap is particularly evident in tabletop robotic manipulation, where memory is essential for solving tasks with partial observability and ensuring robust performance, yet no standardized benchmarks exist. To address this, we introduce MIKASA (Memory-Intensive Skills Assessment Suite for Agents), a comprehensive benchmark for memory RL, with three key contributions: (1) we propose a comprehensive classification framework for memory-intensive RL tasks, (2) we collect MIKASA-Base – a unified benchmark that enables systematic evaluation of memory-enhanced agents across diverse scenarios, and (3) we develop MIKASA-Robo – a novel benchmark of 32 carefully designed memory-intensive tasks that assess memory capabilities in tabletop robotic manipulation. Our contributions establish a unified framework for advancing memory RL research, driving the development of more reliable systems for real-world applications. The code is available at https://sites.google.com/view/memorybenchrobots/.
@@ 1 Introduction
Many real-world problems involve partial observability (Kaelbling et al., 1998), where an agent lacks full access to the environment’s state. These tasks often include sequential decision-making (Chen et al., 2021), delayed or sparse rewards, and long-term information retention (Parisotto et al., 2020; Lampinen et al., 2021). One approach to tackling these challenges is to equip the agent with memory, allowing it to utilize historical information (Meng et al., 2021; Ni et al., 2021).
While there are well-established benchmarks in Natural Language Processing (Bai et al., 2023; An et al., 2023), the evaluation of memory in reinforcement learning (RL) remains fragmented. Existing benchmarks, such as POPGym (Morad et al., 2023), DMLab-30 (Hung et al., 2018) and MemoryGym (Pleines et al., 2023), focus on specific aspects of memory utilization, as they are designed around particular problem domains.
In contrast to classical RL, where benchmarks like Atari (Bellemare et al., 2013) and MuJoCo (Todorov et al., 2012) serve as universal standards, memory-enhanced agents are typically evaluated on custom environments developed alongside their proposals Table 2. This fragmented evaluation landscape obscures important performance variations across different memory tasks. For instance, an agent might excel at maintaining object attributes over extended periods while struggling with sequential recall challenges. Such task-specific strengths and limitations often remain hidden due to narrow evaluation scopes, underscoring the need for a comprehensive benchmark that spans diverse memory-intensive scenarios.
The challenge of memory evaluation becomes particularly evident in robotics. While some robotic tasks naturally involve partial observability, e.g. navigation tasks (Ai et al., 2022; Yadav et al., 2023), many studies artificially create partially observable scenarios from Markov Decision Processes (MDPs) (Kaelbling et al., 1996) by introducing observation noise or masking parts of the state space (Spaan, 2012; Meng et al., 2021; Kurniawati, 2022; Lauri et al., 2023). However, these approaches do not fully capture the complexity of real-world robotic challenges (Lauri et al., 2023), where tasks may require the agent to recall past object configurations, manipulate occluded objects, or perform multi-step procedures that depend heavily on memory.
In this paper, we aim to address these challenges with the following three contributions:
@@
    Memory Tasks Classification: We develop a comprehensive yet practically simple classification of memory-intensive tasks. Our classification framework distills the complex landscape of memory challenges into four essential categories, enabling systematic evaluation while avoiding unnecessary complexity (Figure 1). This approach provides a clear, actionable framework for categorizing and selecting environments that capture fundamental memory challenges in RL and robotics (Section 4).
@@
    Unified Benchmark: We introduce MIKASA-Base, a Gymnasium-based (Towers et al., 2024) framework for evaluating memory-enhanced RL agents (Section 5).
@@
    Robotic Manipulation Tasks: We develop MIKASA-Robo, a suite of 32 carefully designed robotic manipulation tasks that isolate and evaluate specific memory-dependent skills in realistic scenarios (Section 6).
Table 1: MIKASA-Robo: A benchmark comprising 32 memory-intensive robotic manipulation tasks across 12 categories. Each task varies in difficulty and configuration modes. The table specifies episode timeout (T), the necessary information that the agent must memorize in order to succeed (Oracle Info), and task instructions (Prompt) for each environment. See Appendix E for details.
Memory Task Mode Brief description of the task T Oracle Info Prompt Memory Task Type
ShellGame[Mode]-v0 Touch Push Pick Memorize the position of the ball after some time being covered by the cups and then interact with the cup the ball is under. 90 cup_with_ball_number Object
Intercept[Mode]-v0 Slow Medium Fast Memorize the positions of the rolling ball, estimate its velocity through those positions, and then aim the ball at the target. 90 initial_velocity Spatial
InterceptGrab[Mode]-v0 Slow Medium Fast Memorize the positions of the rolling ball, estimate its velocity through those positions, and then catch the ball with the gripper and lift it up. 90 initial_velocity Spatial
RotateLenient[Mode]-v0 Pos PosNeg Memorize the initial position of the peg and rotate it by a given angle. 90 y_angle_diff target_angle Spatial
RotateStrict[Mode]-v0 Pos PosNeg Memorize the initial position of the peg and rotate it to a given angle without shifting its center. 90 y_angle_diff target_angle Object
TakeItBack-v0 Memorize the initial position of the cube, move it to the target region, and then return it to its initial position. 180 xyz_initial Spatial
RememberColor[Mode]-v0 359 Memorize the color of the cube and choose among other colors. 60 true_color_indices Object
RememberShape[Mode]-v0 359 Memorize the shape of the cube and choose among other shapes. 60 true_shape_indices Object
RememberShapeAndColor [Mode]-v0 3$\times$2\3$\times$3 5$\times$3 Memorize the shape and color of the cube and choose among other shapes and colors. 60 true_shapes_info true_colors_info Object
BunchOfColors[Mode]-v0 357 Remember the colors of the set of cubes shown simultaneously in the bunch and touch them in any order. 120 true_color_indices Capacity
SeqOfColors[Mode]-v0 357 Remember the colors of the set of cubes shown sequentially and then select them in any order. 120 true_color_indices Capacity
ChainOfColors[Mode]-v0 357 Remember the colors of the set of cubes shown sequentially and then select them in the same order. 120 true_color_indices Sequential
@@ 2 Related Works
Multiple RL benchmarks are designed to assess agents’ memory capabilities. DMLab-30 (Hung et al., 2018) provides 3D navigation and puzzle tasks, focusing on long-horizon exploration and spatial recall. PsychLab (Leibo et al., 2018) extends DMLab by incorporating tasks that probe cognitive processes, including working memory. MiniGrid and MiniWorld (Chevalier-Boisvert et al., 2023) emphasize partial observability in lightweight 2D and 3D environments, while MiniHack (Samvelyan et al., 2021) builds on NetHack (Küttler et al., 2020), offering small roguelike scenarios that require both short- and long-term memory. BabyAI (Chevalier-Boisvert et al., 2019) combines natural language instructions with grid-based tasks, requiring memory for multi-step command execution. POPGym (Morad et al., 2023) standardizes memory evaluation with tasks ranging from pattern-matching puzzles to complex sequential decision-making. BSuite (Osband et al., 2020) offers a suite of carefully designed experiments that test core RL capabilities, including memory, through controlled tasks on exploration, credit assignment, and scalability. Memory Gym (Pleines et al., 2023) offers a suite of 2D grid environments with partial observability, designed to benchmark memory capabilities in decision-making agents, including endless versions of tasks for evaluating memory over extremely long time intervals. Memory Maze (Pasukonis et al., 2022) presents 3D maze navigation tasks that require memory to solve efficiently.
While these benchmarks offer valuable insights into memory mechanisms, they generally focus on abstract puzzles or navigation tasks. However, none of them fully encompass the broad range of memory utilization scenarios an agent may encounter, and the tasks themselves often differ fundamentally across benchmarks, making direct comparison of memory-enhanced agents difficult.
@@
In the robotics domain, memory requirements become particularly challenging due to the physical nature of manipulation tasks. Unlike abstract environments, robotic manipulation involves complex physical interactions and multi-step procedures that demand both spatial and temporal memory. Existing memory-intensive benchmarks, while useful for diagnostic purposes, struggle to capture these domain-specific challenges. The physical control and object interaction inherent in manipulation tasks introduce additional complexities that are not addressed by traditional memory evaluation frameworks.
@@
Additionally, efforts have been made to classify memory-intensive environments based on specific attributes. For instance, Ni et al. (2023) categorizes these environments into memory/credit assignment, distinguishing them by temporal horizons. Yue et al. (2024) introduces memory dependency pairs, which capture the influence of past events on current decisions, enabling agents to leverage historical context for improved imitation learning in partially observable tasks. Cherepanov et al. (2024a) provides a formal division of agent memory into long-term and short-term depending on the agents’ context length, as well as into declarative and procedural memory depending on the number of environments and episodes, and formalizes the notion of memory-intensive environments. Leibo et al. (2018) takes a different approach by directly adapting established tasks from cognitive psychology and visual psychophysics, providing a standardized way to evaluate agents on well-studied human cognitive benchmarks.
While these classification approaches offer insights into aspects of memory, they overlook physical dimensions in robotics. The interplay between physical interaction and memory remains unexplored, motivating the need for a framework that addresses spatio-temporal aspects in real-world tasks.
@@ 3.1 Partially Observable Markov Decision Process
Partially Observable Markov Decision Process (POMDP) (Kaelbling et al., 1996) extend MDP to account for partial observability, where an agent observes only noisy or incomplete information about the true environments state. POMDP defined by a tuple $(S,A,T,R,\Omega,O,\gamma)$, where:
Table 2: Key memory-intensive environments from the reviewed studies for evaluating agent memory. The Atari (Bellemare et al., 2013) environment with frame stacking is included to illustrate that many memory-enhanced agents are tested solely in MDP. Benchmark first introduced in the same work. Benchmark is open-sourced.
DRQN (Hausknecht & Stone, 2015) DTQN (Esslinger et al., 2022) HCAM (Lampinen et al., 2021) AMAGO (Grigsby et al., 2024) GTrXL (Parisotto et al., 2020) R2I (Samsami et al., 2024) RATE (Cherepanov et al., 2024b) R2A (Goyal et al., 2022) Modified S5 (Lu et al., 2023) Neural Map  (Parisotto & Salakhutdinov, 2017) GBMR (Kang et al., 2024) EMDQN (Lin et al., 2018) MRA (Fortunato et al., 2020) FMRQN (Oh et al., 2016) ADRQN (Zhu et al., 2018) DCEM (Hill et al., 2020) R2D2 (Kapturowski et al., 2018) ERLAM (Zhu et al., 2020a) AdaMemento (Yan et al., 2024) Atari w/o FrameStack Atari with FrameStack gym-gridverse car flag memory card Hallway HeavenHell Ballet Object Permanence DMLab-30 POPGym Passive T-Maze ViZDoom-Two-Colors Numpad Memory Maze Memory Maze (apples) Minigrid-Memory BSuite Goal-Search Doom Maze PsychLab Spot the Difference Goal Navigation Transitive Inference I-Maze Pattern Matching Random Maze Unity Fast-Mapping Task Action Associative Retrieval BabyAI
$S$ is the set of states representing the complete environment configuration; $A$ is the action space; $T(s^{\prime}|s,a):S\times A\times S\to[0,1]$ is the transition function defining the probability of reaching state $s^{\prime}$ from state $s$ after taking action $a$; $R(s,a):S\times A\to\mathbb{R}$ is the reward function specifying the immediate reward for taking action $a$ in state $s$; $\Omega$ is the observation space containing all possible observations; $O(o|s,a):S\times A\times\Omega\to[0,1]$ is the observation function defining the probability of observing $o$ after taking action $a$ and reaching state $s$; $\gamma\in[0,1)$ is the discount factor determining the importance of future rewards. The objective is to find a policy $\pi$ that maximizes the expected discounted cumulative reward: $\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$, where $a_{t}\sim\pi(\cdot|o_{1:t})$ depends on the history of observations rather than the true state. Relying on partial observations makes POMDPs harder to solve than MDPs.
@@ 3.2 Memory-intensive environments
Memory-intensive environment is an environment where agents must leverage past experiences to make decisions, often in problems with long-term dependencies or delayed rewards. More formally, following Cherepanov et al. (2024a), a memory-intensive task is a POMDP where there exists a correlation horizon $\xi>1$, representing the minimum number of timesteps between an event critical for decision-making and when that information must be recalled. Popular memory-intensive environments in RL are listed in Table 2. One way to solving memory-intensive environments is to augment agents with memory mechanisms (see Appendix C).
@@ 3.3 Robotic Tabletop Manipulation
Robotic tabletop manipulation (Shridhar et al., 2022) involves robots manipulating objects on flat surfaces through actions like grasping, pushing, and picking. While crucial for real-world applications (Levine et al., 2018), most existing simulators treat these tasks as MDPs without memory requirements, failing to capture the spatio-temporal dependencies present in real scenarios. This limitation hinders the development of memory-enhanced agents for practical applications.

### v2

# Memory, Benchmark & Robots: A Benchmark for Solving Complex Tasks with Reinforcement Learning

@@ Abstract
Memory is crucial for enabling agents to tackle complex tasks with temporal and spatial dependencies. While many reinforcement learning (RL) algorithms incorporate memory, the field lacks a universal benchmark to assess an agent’s memory capabilities across diverse scenarios. This gap is particularly evident in tabletop robotic manipulation, where memory is essential for solving tasks with partial observability and ensuring robust performance, yet no standardized benchmarks exist. To address this, we introduce MIKASA (Memory-Intensive Skills Assessment Suite for Agents), a comprehensive benchmark for memory RL, with three key contributions: (1) we propose a comprehensive classification framework for memory-intensive RL tasks, (2) we collect MIKASA-Base – a unified benchmark that enables systematic evaluation of memory-enhanced agents across diverse scenarios, and (3) we develop MIKASA-Robo111pip install mikasa-robo-suite – a novel benchmark of 32 carefully designed memory-intensive tasks that assess memory capabilities in tabletop robotic manipulation. Our work introduces a unified framework to advance memory RL research, enabling more robust systems for real-world use. MIKASA is available at https://tinyurl.com/membenchrobots.
@@ 1 Introduction
Many real-world problems involve partial observability [43], where an agent lacks full access to the environment’s state. These tasks often include sequential decision-making [8], delayed or sparse rewards, and long-term information retention [54, 72]. One approach to tackling these challenges is to equip the agent with memory, allowing it to utilize historical information [64, 67]. While there are well-established benchmarks in Natural Language Processing [3, 5], the evaluation of memory in Reinforcement Learning (RL) remains fragmented. Existing benchmarks, such as POPGym [65], DMLab-30 [39] and MemoryGym [75], focus on specific aspects of memory utilization, as they are designed around particular problem domains.
In contrast to classical RL, where benchmarks like Atari [7] and MuJoCo [89] serve as universal standards, memory-enhanced agents are typically evaluated on custom environments developed alongside their proposals Table 2. This fragmented evaluation landscape obscures important performance variations across different memory tasks. For instance, an agent might excel at maintaining object attributes over extended periods while struggling with sequential recall challenges. Such task-specific strengths and limitations often remain hidden due to narrow evaluation scopes, underscoring the need for a comprehensive benchmark that spans diverse memory-intensive scenarios.
The challenge of memory evaluation becomes particularly evident in robotics. While some robotic tasks naturally involve partial observability, e.g. navigation tasks [2, 94], many studies artificially create partially observable scenarios from Markov Decision Processes (MDPs) [42] by introducing observation noise or masking parts of the state space [52, 55, 64, 85]. However, these approaches do not fully capture the complexity of real-world robotic challenges [55], where tasks may require the agent to recall past object configurations, manipulate occluded objects, or perform multi-step procedures that depend heavily on memory. Such tasks include, for example, situations where a service robot needs to memorize occluded objects (e.g., a plate hidden under a towel) or where a home robot needs to accurately wipe the door of a microwave oven several times. Without memory, the robot wouldn’t detect the plate in the first case, and in the second, it would wipe the door endlessly, unsure whether it has cleaned the area or if it’s time to stop.
In this paper, we aim to address these challenges with the following four contributions:
@@
    Memory Tasks Classification. We propose a simple yet comprehensive framework that organizes memory-intensive tasks into four key categories. This structure enables systematic evaluation without added complexity (Figure 1), offering a clear guide for selecting environments that reflect core memory challenges in RL and robotics (Section 4).
@@
    Memory-RL Benchmark. We introduce MIKASA-Base, a Gymnasium-based [90] framework for evaluating memory-enhanced RL agents (Section 5).
@@
    Robotic Manipulation Tasks. We introduce MIKASA-Robo, a suite of 32 robotic tasks targeting specific memory-dependent skills in realistic settings (Section 6), and evaluate them using popular Online RL baselines (Subsection 6.2) and Visual-Language-Action (VLA) models (Subsection 6.4).
  4. 4.
    Robotic Manipulation Datasets. We release datasets for all 32 MIKASA-Robo memory-intensive tasks to support Offline RL research (see Appendix B), and conduct extensive evaluations using a range of Offline RL baselines (Subsection 6.3).
Table 1: MIKASA-Robo: A benchmark comprising 32 memory-intensive robotic manipulation tasks across 12 categories. Each task varies in difficulty and configuration modes. The table specifies episode timeout (T), the necessary information that the agent must memorize in order to succeed (Oracle Info), and task instructions (Prompt) for each environment. See Appendix H for details.
Memory Task Mode Brief description of the task T Oracle Info Prompt Memory
ShellGame Touch Push Pick Memorize the position of the ball after some time being covered by the cups and then interact with the cup the ball is under 90 cup_with_ball_number Object
Intercept Slow Medium Fast Memorize the positions of the rolling ball, estimate its velocity through those positions, and then aim the ball at the target 90 initial_velocity Spatial
InterceptGrab Slow Medium Fast Memorize the positions of the rolling ball, estimate its velocity through those positions, and then catch the ball with the gripper and lift it up 90 initial_velocity Spatial
RotateLenient Pos PosNeg Memorize the initial position of the peg and rotate it by a given angle 90 y_angle_diff target_angle Spatial
RotateStrict Pos PosNeg Memorize the initial position of the peg and rotate it to a given angle without shifting its center 90 y_angle_diff target_angle Spatial
TakeItBack-v0 Memorize the initial position of the cube, move it to the target region, and then return it to its initial position 180 xyz_initial Spatial
RememberColor 3 \ 5 \ 9 Memorize the color of the cube and choose among other colors 60 true_color_indices Object
RememberShape 3 \ 5 \ 9 Memorize the shape of the cube and choose among other shapes 60 true_shape_indices Object
RememberShape- AndColor 3$\times$2 \ 3$\times$3 \ 5$\times$3 Memorize the shape and color of the cube and choose among other shapes and colors 60 true_shapes_info true_colors_info Object
BunchOfColors 3 \ 5 \ 7 Remember the colors of the set of cubes shown simultaneously in the bunch and touch them in any order 120 true_color_indices Capacity
SeqOfColors 3 \ 5 \ 7 Remember the colors of the set of cubes shown sequentially and then select them in any order 120 true_color_indices Capacity
ChainOfColors 3 \ 5 \ 7 Remember the colors of the set of cubes shown sequentially and then select them in the same order 120 true_color_indices Sequential
@@ 2 Related Works
Multiple RL benchmarks are designed to assess agents’ memory capabilities. DMLab-30 [39] provides 3D navigation and puzzle tasks, focusing on long-horizon exploration and spatial recall.
Table 2: Key memory-intensive environments from the reviewed studies for evaluating agent memory. The Atari [7] environment with frame stacking is included to illustrate that many memory-enhanced agents are tested solely in MDP. Benchmark first introduced in the same work. Benchmark is open-sourced.
DRQN [34] DTQN [19] HCAM [54] AMAGO [26] GTrXL [72] R2I [77] RATE [10] R2A [25] Modified S5 [62] Neural Map  [71] GBMR [44] EMDQN [61] MRA [22] FMRQN [69] ADRQN [101] DCEM [36] R2D2 [45] ERLAM [100] AdaMemento [95] Atari w/o FrameStack Atari with FrameStack gym-gridverse car flag memory card Hallway HeavenHell Ballet Object Permanence DMLab-30 POPGym Passive T-Maze ViZDoom-Two-Colors Numpad Memory Maze Memory Maze (apples) Minigrid-Memory BSuite Goal-Search Doom Maze PsychLab Spot the Difference Goal Navigation Transitive Inference I-Maze Pattern Matching Random Maze Unity Fast-Mapping Task Action Associative Retrieval BabyAI
@@
PsychLab [56] extends DMLab by incorporating tasks that probe cognitive processes, including working memory. MiniGrid and MiniWorld [12] emphasize partial observability in lightweight 2D and 3D environments, while MiniHack [78] builds on NetHack [53], offering small roguelike scenarios that require both short- and long-term memory. BabyAI [11] combines natural language instructions with grid-based tasks, requiring memory for multi-step command execution. POPGym [65] standardizes memory evaluation with tasks ranging from pattern-matching puzzles to complex sequential decision-making. BSuite [70] offers a suite of carefully designed experiments that test core RL capabilities, including memory, through controlled tasks on exploration, credit assignment, and scalability. Memory Gym [75] offers a suite of 2D grid environments with partial observability, designed to benchmark memory capabilities in decision-making agents, including endless versions of tasks for evaluating memory over extremely long time intervals. Memory Maze [73] presents 3D maze navigation tasks that require memory to solve efficiently.
@@
While these benchmarks offer valuable insights into memory mechanisms, they generally focus on abstract puzzles or navigation tasks. However, none of them fully encompass the broad range of memory utilization scenarios an agent may encounter, and the tasks themselves often differ fundamentally across benchmarks, making direct comparison of memory-enhanced agents difficult. In the robotics domain, memory requirements become particularly challenging due to the physical nature of manipulation tasks. Unlike abstract environments, robotic manipulation involves complex physical interactions and multi-step procedures demanding both spatial and temporal memory. Existing memory-intensive benchmarks, while useful for diagnostic purposes, struggle to capture these domain-specific challenges. The physical control and object interaction inherent in manipulation tasks introduce additional complexities not addressed by traditional memory evaluation frameworks.
@@
Efforts have been made to classify memory-intensive environments by specific attributes. For example, Ni et al. [68] divides them into memory/credit assignment based on temporal horizons. Yue et al. [97] proposes memory dependency pairs to model how past events influence current decisions, aiding imitation learning in partially observable tasks. Cherepanov et al. [9] defines agent memory types: long-term vs. short-term (based on context length), and declarative vs. procedural (based on environments and episodes), and formalizes memory-intensive environments. Leibo et al. [56] instead adapts tasks from cognitive psychology and psychophysics to evaluate agents on human cognitive benchmarks. While these classifications highlight aspects of memory, they overlook physical dimensions in robotics. The link between physical interaction and memory remains underexplored, motivating a framework for spatio-temporal memory in real-world tasks.
Concurrent with our work Fang et al. [21] also proposed MemoryBench, a benchmark for memory-intensive manipulation consisting of only three tasks designed to access only one type of memory, spatial memory. This benchmark is based on RLBench [40], which does not allow efficient parallelization of training.
@@ 3.1 Partially Observable Markov Decision Process
Partially Observable Markov Decision Process (POMDP) [42] extend MDP to account for partial observability, where an agent observes only noisy or incomplete information about the true environments state. POMDP defined by a tuple $(S,A,T,R,\Omega,O,\gamma)$, where: $S$ is the set of states representing the complete environment configuration; $A$ is the action space; $T(s^{\prime}|s,a):S\times A\times S\to[0,1]$ is the transition function defining the probability of reaching state $s^{\prime}$ from state $s$ after taking action $a$; $R(s,a):S\times A\to\mathbb{R}$ is the reward function specifying the immediate reward for taking action $a$ in state $s$; $\Omega$ is the observation space containing all possible observations; $O(o|s,a):S\times A\times\Omega\to[0,1]$ is the observation function defining the probability of observing $o$ after taking action $a$ and reaching state $s$; $\gamma\in[0,1)$ is the discount factor determining the importance of future rewards. The objective is to find a policy $\pi$ that maximizes the expected discounted cumulative reward: $\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$, where $a_{t}\sim\pi(\cdot|o_{1:t})$ depends on the history of observations rather than the true state. Relying on partial observations makes POMDPs harder to solve than MDPs.
@@ 3.2 Memory-intensive environments
Memory-intensive environment is an environment where agents must leverage past experiences to make decisions, often in problems with long-term dependencies or delayed rewards. More formally, following [9], a memory-intensive task is a POMDP where there exists a correlation horizon $\xi>1$, representing the minimum number of timesteps between an event critical for decision-making and when that information must be recalled. Popular memory-intensive environments in RL are listed in Table 2. One way to solving memory-intensive environments is to augment agents with memory mechanisms (see Appendix E).
@@ 3.3 Robotic Tabletop Manipulation
Robotic tabletop manipulation [80] involves robots manipulating objects on flat surfaces through actions like grasping, pushing, and picking. While crucial for real-world applications [57], most existing simulators treat these tasks as MDPs without memory requirements, failing to capture the spatio-temporal dependencies present in real scenarios. This limitation hinders the development of memory-enhanced agents for practical applications.